OpenStudy (mendicant_bias):

Doing another integrodifferential equation now:

OpenStudy (mendicant_bias):

\[f(t)+\int\limits_{0}^{t}f(\tau)d \tau = 1\]

OpenStudy (mendicant_bias):

Making the identification that this is the same sort of situation as \[\int\limits_{0}^{t}f(\tau)d \tau = \mathcal{L^{-1}}\left\{\vphantom{}\frac{F(s)}{s}\right\}\]

OpenStudy (mendicant_bias):

Alright, uhh, Oh! This is really odd, can I do this? Let's see...

OpenStudy (sidsiddhartha):

yes this is easier than the first one

OpenStudy (mendicant_bias):

LAtEx Isn't playing nice (whoops lol), but I think we're on the same page, inverse and forward operation cancel each other out

ganeshie8 (ganeshie8):

you can think of \(\int\limits_0^t f(\tau)d\tau \) as \(f(t)*1\) right ? why to identiy it as some other situation ?

OpenStudy (mendicant_bias):

ughh gonna rewrite stuff, one moment \[F(s)+\mathcal{L}\left\{\mathcal{L^{-1}}\left\{\vphantom{}\frac{F(s)}{s}\right\} \right\}=1\]

OpenStudy (mendicant_bias):

\[F(s)+\frac{F(s)}{s}=1\]

ganeshie8 (ganeshie8):

you need to transform right side also

OpenStudy (sidsiddhartha):

yes

OpenStudy (sidsiddhartha):

forgot that 1

OpenStudy (mendicant_bias):

Whoops! Thank you, heh\[F(s)\bigg[1+\frac{1}{s}\bigg]=\frac{1}{s}\]

OpenStudy (sidsiddhartha):

yeah that fine now just inverse

OpenStudy (mendicant_bias):

\[F(s)=\frac{1+s}{s}\](Trying to go quickly, might've jsut made an algebra error)

OpenStudy (sidsiddhartha):

yes lol

OpenStudy (sidsiddhartha):

\[F(s)\frac{ 1+s }{ s }=1/s\]

OpenStudy (sidsiddhartha):

\[F(s)=\frac{ 1 }{ s+1 }\]

ganeshie8 (ganeshie8):

i would multiply \(s\) through out so that fractions disappear as i hate fractions

OpenStudy (mendicant_bias):

Ah, I see what you did, yeah, makes it way easier

OpenStudy (mendicant_bias):

\[\mathcal{L^{-1}}\left\{\vphantom{}\frac{1}{s+1}\right\}=e^{-t}\]

OpenStudy (sidsiddhartha):

woo-hoo

OpenStudy (mendicant_bias):

Alright, ~4 hours left

OpenStudy (mendicant_bias):

https://www.youtube.com/watch?v=oOM-91dydPU

OpenStudy (mendicant_bias):

Time to start a new problem.

ganeshie8 (ganeshie8):

I can see you're ready !

ganeshie8 (ganeshie8):

*for the exam

OpenStudy (mendicant_bias):

Heh, trying my best. We'll see, this professor is pretty tough.

OpenStudy (sidsiddhartha):

ohh i like that one :)